OpenStudy (anonymous):

what is the slope of the line perpendicular to y=2x-8

Vocaloid (vocaloid):

find the slope and take the negative reciprocal

Vocaloid (vocaloid):

what's the slope of y = 2x - 8?

OpenStudy (anonymous):

Slope = 4.000/2.000 = 2.000 x-intercept = 8/2 = 4 y-intercept = -8/1 = -8.00000